OpenStudy (anonymous):

hello, the problem i am trying to solve says to find the exact value of each of the trigonometric functions theta. the one it gives me is cosine of theta equals 4/7, 270degrees

OpenStudy (anonymous):

draw a triangle

OpenStudy (anonymous):



OpenStudy (anonymous):

all you need is \(y\) which, by pythagoras, is \[\sqrt{7^2-4^2}=\sqrt{33}\] now you can take any trig ratio you like just keep in mind you are in quadrant 4 so be careful with the signs

OpenStudy (anonymous):

could you help me find the first one please? i have no idea what steps need to be done and what an acceptable answer would look like, just so i can get the hang of it.